For a science experiment you need to electroplate zinc onto both sides of a very thin, 1.00cm
Answer #1

determine the number of zinc atoms in the plating, multiply by 2 to get the number of electrons needed, convert to Coulombs and divide by the current.
atomic weight of zinc Wz = 65.38 g/mole
Zinc atom mass mz = Wz/(1000*Avogadro) = 1.0857E-25 kg
Zinc atom volume Vz = mz/rho = 1.5206E-29 m^3
V(plating) = 150E-9*2*0.01^2 = 2.8E-11 m^3
Natoms = V(plating)/Vz = 1.841E18
Nelectrons = 2*Natoms = 3.683E18
Total charge Q = Nelectrons*1.6022E-19 = 0.59009026 C
t = Q/I = 421.493 s = 421.5 sec appro